### v1

@@ 2 A relative-type inexact pALM
with $\Delta_{x}^{k+1}:=\bm{x}^{k+1}-\bm{x}^{k}$ and $\Delta_{y}^{k+1}:=\bm{y}^{k+1}-\bm{y}^{k}$, to guarantee an asymptotically Q-(super)linear convergence rate. [^2] Note that the error criterion (A) in (2.7) alone is sufficient for establishing the global convergence of the Snipal; see [21, Section 3]. Both error criteria (A) and (B) in (2.7) are of the absolute type, meaning that they require the pre-specification of two summable sequences of tolerance parameters, $\{\varepsilon_{k}\}\subseteq[0,\infty)$ and $\{\delta_{k}\}\subseteq[0,1)$, to control the error incurred in the inexact minimization of the subproblem. Since there is generally no direct guidance on optimally selecting these tolerance parameters to achieve good convergence efficiency, this absolute-type criterion typically requires careful tuning for the tolerance parameters, which may make Snipal less user-friendly in practical implementations. In contrast, the error criterion (2.5) used in ripALM is of the relative type, meaning that the error $2\big{|}\langle\bm{w}^{k}-\bm{y}^{k+1},\,\sigma_{k}\Delta^{k+1}\rangle\big{|}+\big{\|}\sigma_{k}\Delta^{k+1}\big{\|}^{2}$ is regulated by a tentative successive difference related to the progress of the algorithm. Moreover, this relative-type criterion only involves a single tolerance parameter $\rho\in[0,1)$, thus simplifying the process of parameter tuning both computationally and in terms of implementation, as we shall see in subsection 5.1.
@@
Thanks to the advantage of eliminating the need to select an infinite sequence of tolerance parameters, various versions of relative error criteria have been widely adopted in numerous well-known algorithms (e.g., PPA, ALM, ADMM, FISTA, etc) to approximately solve subproblems over the past two decades. This trend began with the seminal works of Solodov and Svaiter [42, 43, 44, 45, 46], which has since inspired the development of numerous algorithms; see, for example, [3, 11, 12, 30, 49, 51, 52]. Following the same research theme, Yang et al. [51] recently developed a corrected inexact proximal augmented Lagrangian method (ciPALM) for solving a class of group-quadratic regularized optimal transport problems. The error criterion used there can be described as follows: choose $\rho\in[0,1)$, at each iteration, approximately solve the subproblem (2.4) to find a point $\widetilde{\bm{y}}^{k+1}$ such that
$\widetilde{\Delta}^{k+1}:=$ $~{}\nabla_{y}\mathcal{L}_{\sigma_{k}}(\widetilde{\bm{y}}^{k+1},\bm{x}^{k})+\tau_{k}\sigma_{k}^{-1}(\widetilde{\bm{y}}^{k+1}-\bm{y}^{k}),$ (2.8)
@@
$\big{\|}{\sigma_{k}}\widetilde{\Delta}^{k+1}\big{\|}^{2}\leq$ $~{}\rho\min\left({\tau_{k}},\,1\right)\left\|\texttt{prox}_{\sigma_{k}f}\big{(}\bm{x}^{k}+\sigma_{k}{A}^{\top}\widetilde{\bm{y}}^{k+1}\big{)}-\bm{x}^{k}\right\|^{2}+\tau_{k}\left\|\widetilde{\bm{y}}^{k+1}-\bm{y}^{k}\right\|^{2}.$
@@
When (2.8) is satisfied, the multiplier is updated as usual: $\bm{x}^{k+1}=\texttt{prox}_{\sigma_{k}f}\big{(}\bm{x}^{k}+\sigma_{k}{A}^{\top}\widetilde{\bm{y}}^{k+1}\big{)}$. This is then followed by an extra correction step: $\bm{y}^{k+1}=\bm{y}^{k}-\tau_{k}^{-1}\sigma_{k}\left(A\bm{x}^{k+1}-\bm{b}\right)$. After these updates, the algorithm proceeds to the next iteration. One can see that the error criterion (2.8) used in the ciPALM is also of the relative type and looks even simpler than (2.5) used in our ripALM. However, we should point out that the ciPALM requires an extra correction step to guarantee the convergence, and hence it may only be viewed as a pALM-like algorithm. In contrast, our relative error criterion (2.5) is designed for the vanilla pALM, where the error variable $\bm{w}^{k}$ is used solely in the construction of the error criterion (2.5) and does not directly influence either the objective function in (2.4) or the updating rules of the primal and dual variables. Our experimental results also show that the correction step tends to slow down the convergence of the pALM, whereas our ripALM can offer greater robustness and efficiency; see subsection 5.1 for numerical comparisons.
@@
Our relative error criterion (2.5) is inspired by Eckstein and Silva’s practical relative error criterion [11], which was developed for the approximate minimization of the subproblems in the vanilla ALM (i.e., without the proximal term $\frac{\tau_{k}}{2\sigma_{k}}\big{\|}\bm{y}-\bm{y}^{k}\big{\|}^{2}$ in the subproblem (2.4) in our context). Unlike their work, we suggest incorporating $\frac{\tau_{k}}{2\sigma_{k}}\big{\|}\bm{y}-\bm{y}^{k}\big{\|}^{2}$ in the subproblem (2.4). This proximal term guarantees the existence and uniqueness of the optimal solution of the strongly convex subproblem (2.4). Introducing the proximal term also ensures the positive definiteness of the generalized Hessian of the objective function in (2.4), thereby facilitating the application of the semi-smooth Newton method to solve the subproblem (2.4) effectively, as shown in Section 4. In addition, as we will see later in the next section, our ripALM not only shares the same convergence properties for the sequence $\{\bm{x}^{k}\}$ as Eckstein and Silva’s relative-type inexact ALM developed in [1, 11, 54], but also offers other theoretical advantages in that the sequence $\{\bm{y}^{k}\}$ is guaranteed to converge, and both primal and dual sequences achieve asymptotically Q-(super)linear convergence rates under a relatively weaker error bound assumption.
@@ 3 Convergence analysis
In this section, we study the convergence properties of ripALM in Algorithm 1. Before proceeding, we recall the definition of $G$ from (2.2) and further define $F:\mathbb{R}^{N}\times\mathbb{R}^{M}\rightrightarrows\mathbb{R}^{N}\times\mathbb{R}^{M}$ to be the concave conjugate of $G$:
@@
$\partial\ell(\bm{y},\bm{x}):=\left\{A\bm{x}-\bm{b}\right\}\times\left\{v-A^{\top}\bm{y}\mid v\in\partial f(\bm{x})\right\}.$
Moreover, all the subgradient maps $\partial G$, $\partial F$ and $\partial\ell$ are maximal monotone operators, and satisfy that (see also [11, equation (23)]):
@@
If $(\bm{y}^{*},\bm{x}^{*})\in\mathbb{R}^{M}\times\mathbb{R}^{N}$ satisfies $(\bm{0},\bm{0})\in\partial\ell(\bm{y}^{*},\bm{x}^{*})$, then $\bm{y}^{*}$ solves problem (2.1) (i.e., problem (1.2)) and $\bm{x}^{*}$ solves problem (3.2) (i.e., problem (1.1)). In this case, we call $(\bm{y}^{*},\bm{x}^{*})$ a saddle point of the Lagrangian function $\ell(\bm{y},\bm{x})$. If such a saddle point exists, then strong duality holds, that is, the optimal values of problems (2.1) and (3.2) are finite and equal, i.e., $G(\bm{y}^{*},\bm{0})=F(\bm{0},\bm{x}^{*})$. Moreover, the set of saddle points can be written as $\mathcal{Y}^{*}\times\mathcal{X}^{*}\subseteq\mathbb{R}^{M}\times\mathbb{R}^{N}$, where $\mathcal{Y}^{*}$ is the solution set of problem (2.1) (i.e., problem (1.2)) and $\mathcal{X}^{*}$ is the solution set of problem (3.2) (i.e., problem (1.1)).
With the above preparations, we are now ready to establish the convergence results of the proposed ripALM in Algorithm 1. The analysis is inspired by [1, 11, 54], but is more involved. We need to be particularly careful in establishing a proper recursive relation with respect to $\bm{x}^{k}$, $\bm{y}^{k}$ and $\bm{w}^{k}$, due to the inclusion of the proximal term $\frac{\tau_{k}}{2\sigma_{k}}\|\bm{y}-\bm{y}^{k}\|^{2}$.
@@
[…]
@@
    Both the sequences $\left\{G(\bm{y}^{k+1},\bm{\xi}^{k+1})\right\}$ and $\left\{F(\bm{\theta}^{k+1},\,\bm{x}^{k+1})\right\}$ converge to the common optimal value of problems (2.1) and (3.2).
@@
We next study the asymptotically Q-(super)linear convergence rate of our ripALM under an error-bound condition presented in Assumption A. As noted in [21, Lemma 2.4], this error-bound condition is weaker than the local upper Lipschitz continuity of $(\partial\ell)^{-1}$ at the origin. The latter condition was used in [54] to establish the asymptotic Q-(super)linear convergence rate for Eckstein and Silva’s relative-type inexact ALM, while the former, weaker condition has also been employed in [21] and [51] to establish the asymptotic Q-(super)linear convergence rate for the Snipal and ciPALM, respectively.
@@
For any $r>0$, there exist a constant $\kappa>0$ such that, for any $(\bm{y},\bm{x})\in\left\{(\bm{y},\bm{x})\in\mathbb{R}^{M}\times\mathbb{R}^{N}\mid\mathrm{dist}\left((\bm{y},\bm{x}),\,(\partial\ell)^{-1}(\bm{0},\bm{0})\right)\leq r\right\}$,
@@
Under the same assumptions as in Theorem 3.1, suppose additionally that Assumption A holds, and the sequences of parameters $\rho$, $\{\sigma_{k}\}$ and $\{\tau_{k}\}$ satisfy that
$\sqrt{\tau_{\min}}-2\sqrt{\rho}>0\quad\text{and}\quad\liminf\limits_{k\to\infty}~{}\sigma_{k}>c\cdot\frac{2\kappa\sqrt{\tau_{\max}}\left(\rho+\sqrt{\rho\max\left\{1,\tau_{\max}\right\}}\right)}{\sqrt{\tau_{\min}}-2\sqrt{\rho}},$ (3.5)
where $c>1$ is an arbitrarily given positive constant and $\tau_{\max}:=\tau_{0}\prod_{k=0}^{\infty}(1+\nu_{k})$. Let $\Lambda^{k}:=\operatorname{Diag}(\tau_{k}I_{M},I_{N})$, $\overline{\tau}_{k}:=\max\left\{1,\tau_{k}\right\}$ and
@@
    For any sufficiently large $k$, we have that $\gamma_{k}>0$ and
@@
    $\limsup\limits_{k\to\infty}\,\left\{\mu_{k}:=\sqrt{\frac{1+\nu_{k}}{1+\gamma_{k}}}\right\}<1.$
@@
Remark 3.1 (Comments on $\mu_{k}$).
One can see from the expression of $\gamma_{k}$ that, after a finite number of iterations, $\gamma_{k}$ becomes proportional to the penalty parameter $\sigma_{k}$, provided that $\rho$, $\{\sigma_{k}\}$ and $\{\tau_{k}\}$ satisfy conditions in (3.5). Therefore, if $\sqrt{\tau_{\min}}-2\sqrt{\rho}>0$, the coefficient $\mu_{k}$ can be less than 1 when $\sigma_{k}$ is sufficiently large, and it approaches 0 as $\sigma_{k}$ tends to +$\infty$+ ∞. This readily demonstrates that the sequence $\left\{(\bm{y}^{k},\bm{x}^{k})\right\}$ converges to the set of saddle points at a Q-(super)linear rate if $\sigma_{k}$ is sufficiently large and $\sqrt{\tau_{\min}}-2\sqrt{\rho}>0$. In practical implementations, one could simply choose an increasing sequence of $\{\sigma_{k}\}$ with $\sigma_{k}\uparrow+\infty$, and set $\tau_{\min}\geq 4$ to ensure that $\sqrt{\tau_{\min}}-2\sqrt{\rho}>0$ for any $\rho\in[0,1)$. In contrast, as discussed in [51, Remark 2.1], the ciPALM (which is also of relative-type but uses the error criterion (2.8)) should require $\rho<\frac{1}{3}$ to guarantee an asymptotically Q-(super)linear convergence rate under the same error bound condition. This highlights another advantage of our ripALM, as it offers greater flexibility in choosing $\rho$.
4 Application to the QROT problem
In this section, we will present the application of ripALM in Algorithm 1 for solving quadratically regularized optimal transport (QROT) problems. As a significant variant of the classical optimal transport problem, the QROT problem has garnered notable attention in recent years; see, e.g., [4, 13, 28] for more details. Mathematically, the QROT problem is given as follows:
$\min\limits_{X}~{}\Big{\{}\frac{\lambda}{2}\|X\|_{F}^{2}+\langle C,\,X\rangle\;\mid\;X\bm{1}_{n}=\bm{\alpha},~{}~{}X^{\top}\bm{1}_{m}=\bm{\beta},~{}~{}X\geq 0\Big{\}},$ (4.1)

### v2

@@ 2 A relative-type inexact pALM
with $\Delta_{x}^{k+1}:=\bm{x}^{k+1}-\bm{x}^{k}$ and $\Delta_{y}^{k+1}:=\bm{y}^{k+1}-\bm{y}^{k}$, to guarantee an asymptotic (super)linear convergence rate. [^3] Note that the error criterion (A) in (2.7) alone is sufficient for establishing the global convergence of the Snipal; see [24, Section 3]. Both error criteria (A) and (B) in (2.7) are of the absolute type, meaning that they require the pre-specification of two summable tolerance parameter sequences, $\{\varepsilon_{k}\}\subseteq[0,\infty)$ and $\{\delta_{k}\}\subseteq[0,1)$, to control the error incurred in the inexact minimization of the subproblem. Since there is generally no direct guidance on optimally selecting these tolerance parameters to achieve good convergence efficiency, this absolute-type criterion typically requires careful tuning for the tolerance parameters, which may make Snipal less user-friendly in practical implementations. In contrast, the error criterion (2.5) used in ripALM is of the relative type, meaning that the error $2\big|\langle\bm{w}^{k}-\bm{y}^{k+1},\,\sigma_{k}\Delta^{k+1}\rangle\big|+\big\|\sigma_{k}\Delta^{k+1}\big\|^{2}$ is regulated by a tentative successive difference related to the progress of the algorithm. Moreover, this relative-type criterion only involves a single tolerance parameter $\rho\in[0,1)$, thus simplifying the process of parameter tuning in both computation and implementation, as we shall see in subsection 5.1.
@@
Thanks to the advantage of eliminating the need to select an infinite sequence of tolerance parameters, relative-type error criteria have been widely adopted in numerous well-known algorithms (e.g., PPA, ALM, ADMM, FISTA, etc.) to approximately solve subproblems over the past two decades. This trend began with the seminal works of Solodov and Svaiter [45, 46, 47, 48, 49], which has since inspired the development of numerous relative-type inexact algorithms; see, for example, [4, 14, 15, 33, 56, 58, 59]. Following this line of research, Yang et al. [58] recently developed a corrected inexact proximal augmented Lagrangian method (ciPALM) for solving a class of group-quadratic regularized optimal transport problems. The error criterion used there can be described as follows: choose $\rho\in[0,1)$, at each iteration, approximately solve the subproblem (2.4) to find a point $\widetilde{\bm{y}}^{k+1}$ such that
$\widetilde{\Delta}^{k+1}$= $~\nabla_{y}\mathcal{L}_{\sigma_{k}}(\widetilde{\bm{y}}^{k+1},\bm{x}^{k})+\tau_{k}\sigma_{k}^{-1}(\widetilde{\bm{y}}^{k+1}-\bm{y}^{k}),$ (2.8)
@@
$\hskip-7.11317pt\big\|{\sigma_{k}}\widetilde{\Delta}^{k+1}\big\|^{2}\leq$ $~\rho^{2}\min\left({\tau_{k}},\,1\right)\left(\left\|\mathtt{prox}_{\sigma_{k}f}\big(\bm{x}^{k}+\sigma_{k}{A}^{\top}\widetilde{\bm{y}}^{k+1}\big)-\bm{x}^{k}\right\|^{2}+\tau_{k}\left\|\widetilde{\bm{y}}^{k+1}-\bm{y}^{k}\right\|^{2}\right).$
@@
When (2.8) is satisfied, the multiplier is updated as usual: $\bm{x}^{k+1}=\mathtt{prox}_{\sigma_{k}f}\big(\bm{x}^{k}+\sigma_{k}{A}^{\top}\widetilde{\bm{y}}^{k+1}\big)$. This is then followed by an extra correction step: $\bm{y}^{k+1}=\bm{y}^{k}-\tau_{k}^{-1}\sigma_{k}\left(A\bm{x}^{k+1}-\bm{b}\right)$. After these updates, the algorithm proceeds to the next iteration. One can see that the error criterion (2.8) used in ciPALM is also of the relative type and looks even simpler than (2.5) used in our ripALM. However, we should point out that ciPALM requires an additional correction step to ensure its convergence, and hence it may only be viewed as a pALM-like algorithm. Moreover, our numerical results indicate that this correction step tends to degrade the practical performance of the pALM framework. This raises the natural question of whether one can design a relative-type error criterion directly for the vanilla pALM framework. Addressing this question serves as the main motivation of this work. To this end, we propose the relative error criterion (2.5), where the error variable $\bm{w}^{k}$ is used solely in the construction of the error criterion (2.5) and does not directly influence either the objective function in (2.4) or the updating rules of the primal and dual variables. Our numerical results show that the proposed ripALM may offer greater robustness and efficiency; see subsection 5.1 for numerical comparisons.
@@
Our relative error criterion (2.5) is inspired by Eckstein and Silva’s practical relative error criterion [14], which was developed for the approximate minimization of the subproblems in the vanilla ALM (i.e., without the proximal term $\frac{\tau_{k}}{2\sigma_{k}}\big\|\bm{y}-\bm{y}^{k}\big\|^{2}$ in the subproblem (2.4) in our context). Unlike their work, we suggest incorporating $\frac{\tau_{k}}{2\sigma_{k}}\big\|\bm{y}-\bm{y}^{k}\big\|^{2}$ in the subproblem (2.4). This proximal term guarantees the existence and uniqueness of the optimal solution of the strongly convex subproblem (2.4). Introducing the proximal term also ensures the positive definiteness of the generalized Hessian/Jacobian of the objective function in (2.4), thereby facilitating the application of the semi-smooth Newton method to the subproblem (2.4) effectively, as shown in Section 4.1. In addition, as we will see later in the next section, our ripALM not only shares the same convergence properties for the sequence $\{\bm{x}^{k}\}$ as Eckstein and Silva’s relative-type inexact ALM (denoted by ES-iALM) developed in [1, 14, 61], but also offers additional theoretical advantages: the sequence $\{\bm{y}^{k}\}$ is guaranteed to be bounded; moreover, under a relatively weaker error bound assumption (than the local upper Lipschitz continuity of $(\partial\ell)^{-1}$ at the origin), the sequence $\{\bm{y}^{k}\}$ is further proved to be convergent, and both primal and dual sequences achieve asymptotic (super)linear convergence rates.
@@ 3 Convergence analysis
In this section, we study the convergence properties of the ripALM in Algorithm 1. Before proceeding, we recall the definition of $G$ from (2.2) and further define $F:\mathbb{R}^{M}\times\mathbb{R}^{N}\rightrightarrows\mathbb{R}^{M}\times\mathbb{R}^{N}$ to be the concave conjugate of $G$:
@@
$\partial\ell(\bm{y},\bm{x})=\left\{A\bm{x}-\bm{b}\right\}\times\left\{\bm{v}-A^{\top}\bm{y}\mid\bm{v}\in\partial f(\bm{x})\right\}.$
Moreover, all the subgradient maps $\partial G$, $\partial F$, and $\partial\ell$ are maximal monotone operators, and satisfy that (see also [14, equation (23)]):
@@
If $(\bm{y}^{*},\bm{x}^{*})\in\mathbb{R}^{M}\times\mathbb{R}^{N}$ satisfies $(\bm{0},\bm{0})\in\partial\ell(\bm{y}^{*},\bm{x}^{*})$, i.e, $(\bm{y}^{*},\bm{x}^{*})$ solves the following KKT system:
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\bm{0}\in\partial f(\bm{x})-A^{\top}\bm{y},\quad A\bm{x}-\bm{b}=\bm{0},$
then $\bm{y}^{*}$ solves problem (2.1) (i.e., problem (1.2)) and $\bm{x}^{*}$ solves problem (3.2) (i.e., problem (1.1)). In this case, we call $(\bm{y}^{*},\bm{x}^{*})$ a saddle point of the Lagrangian function $\ell(\bm{y},\bm{x})$. If such a saddle point exists, then the strong duality holds, that is, the optimal values of problems (2.1) and (3.2) are finite and equal, i.e., $G(\bm{y}^{*},\bm{0})=F(\bm{0},\bm{x}^{*})$. Moreover, the set of saddle points can be written as $\mathcal{Y}^{*}\times\mathcal{X}^{*}\subseteq\mathbb{R}^{M}\times\mathbb{R}^{N}$, where $\mathcal{Y}^{*}$ is the solution set of problem (2.1) (i.e., problem (1.2)) and $\mathcal{X}^{*}$ is the solution set of problem (3.2) (i.e., problem (1.1)).
With the above preparations, we are now ready to establish the convergence results of the proposed ripALM in Algorithm 1. Note that, while our analysis is inspired by that of ES-iALM in [1, 14, 61], the inclusion of the proximal term $\frac{\tau_{k}}{2\sigma_{k}}\|\bm{y}-\bm{y}^{k}\|^{2}$ indeed introduces new technical challenges. We need to carefully derive new recursive inequalities that simultaneously involve both primal and dual sequences, $\{\bm{x}^{k}\}$ and $\{\bm{y}^{k}\}$, which cannot be obtained by applying existing results directly. In particular, to establish the asymptotic (super)linear convergence rate in Theorem 3.2, we derive a novel recursion (A.20), which critically relies on the presence of a strictly positive proximal parameter $\tau_{k}$, as $\tau_{k}$ appears in the denominator of the coefficient on the right-hand side. Consequently, this novel recursion is tailored to our ripALM framework.
@@
Let the functions $G$, $F$ and $\ell$ be defined as in (2.2), (3.1) and (2.3), respectively. Let $\rho\in[0,1)$, $\{\sigma_{k}\}$ be a positive sequence satisfying that $\sigma_{k}\geq\sigma_{\min}>0$ for all $k\geq 0$, and $\{\tau_{k}\}$ be a positive sequence satisfying that
@@
    Both sequences $\left\{G(\bm{y}^{k+1},\bm{\xi}^{k+1})\right\}$ and $\left\{F(\bm{\theta}^{k+1},\,\bm{x}^{k+1})\right\}$ converge to the common optimal value of problems (2.1) and (3.2).
@@
Next, we study the asymptotic (super)linear convergence rate of our ripALM under an error bound condition presented in Assumption A. As noted in [24, Lemma 2.4], this error-bound condition is weaker than the local upper Lipschitz continuity of $(\partial\ell)^{-1}$ at the origin. The latter condition was used in [61] to establish the asymptotic (super)linear convergence rate for Eckstein and Silva’s relative-type inexact ALM, while the former, weaker condition has also been employed in [24] and [58] to establish the asymptotic (super)linear convergence rate for Snipal and ciPALM, respectively.
@@
For any $r>0$, there exists a constant $\kappa>0$ such that, for any $(\bm{y},\bm{x})\in\left\{(\bm{y},\bm{x})\in\mathbb{R}^{M}\times\mathbb{R}^{N}\mid\mathrm{dist}\left((\bm{y},\bm{x}),\,(\partial\ell)^{-1}(\bm{0},\bm{0})\right)\leq r\right\}$,
@@
Let $\ell$ be defined as in (2.3), and let $\rho\in[0,1)$, $\{\sigma_{k}\}$ be a positive sequence satisfying that $\sigma_{k}\geq\sigma_{\min}>0$ for all $k\geq 0$, and $\{\tau_{k}\}$ be a positive sequence satisfying that
Suppose additionally that $\ell$ admits a saddle point (i.e., $(\partial\ell)^{-1}(\bm{0},\bm{0})\neq\emptyset$), Assumption A holds, and the sequences of parameters $\rho$, $\{\sigma_{k}\}$ and $\{\tau_{k}\}$ satisfy that
$\sqrt{\tau_{\min}}-2\sqrt{\rho}>0\quad\text{and}\quad\liminf\limits_{k\to\infty}~\sigma_{k}>c\cdot\frac{2\kappa\sqrt{\tau_{\max}}\left(\rho+\sqrt{\rho\,\overline{\tau}_{\max}}\right)}{\sqrt{\tau_{\min}}-2\sqrt{\rho}},$ (3.5)
where $c>1$ is an arbitrarily given positive constant, $\tau_{\max}:=\tau_{0}\prod_{k=0}^{\infty}(1+\nu_{k})$, and $\overline{\tau}_{\max}:=\max\left\{1,\tau_{\max}\right\}$. Let $\Lambda^{k}:=\operatorname{Diag}(\tau_{k}I_{M},I_{N})$, $\overline{\tau}_{k}:=\max\left\{1,\tau_{k}\right\}$ and
@@
    For all sufficiently large $k$, we have that
    $\gamma_{k}\geq\left(\frac{c-1}{c}\right)\cdot\frac{\sqrt{\tau_{\min}}-2\sqrt{\rho}}{\sqrt{\tau_{\min}}}\cdot\frac{\sigma_{k}^{2}}{\kappa^{2}\left(\sqrt{\rho}+\sqrt{\overline{\tau}_{\max}}\right)^{2}\overline{\tau}_{\max}}>0,$
    and
@@
    $\mu_{k}:=\sqrt{\frac{1+\nu_{k}}{1+\gamma_{k}}}\quad\mbox{satisfies}\quad\limsup\limits_{k\to\infty}\,\left\{\mu_{k}\right\}<1~~\mbox{as}~~\nu_{k}\to 0.$
@@
Remark 3.1 (Comments on $\gamma_{k}$ and $\mu_{k}$).
One can see from the expression of $\gamma_{k}$ in Theorem 3.2 that, after a finite number of iterations, $\gamma_{k}$ becomes proportional to the squared penalty parameter $\sigma_{k}^{2}$, provided that $\rho$, $\{\sigma_{k}\}$ and $\{\tau_{k}\}$ satisfy the conditions in (3.5). Therefore, if $\sqrt{\tau_{\min}}-2\sqrt{\rho}>0$, the convergence factor $\mu_{k}$ can be less than 1 when $\sigma_{k}$ is sufficiently large, and it approaches 0 as $\sigma_{k}$ tends to $\infty$. This readily demonstrates that the sequence $\left\{(\bm{y}^{k},\bm{x}^{k})\right\}$ converges to the set of saddle points at a Q-(super)linear rate if $\sigma_{k}$ is sufficiently large and $\sqrt{\tau_{\min}}-2\sqrt{\rho}>0$. In practical implementations, one could simply choose an increasing sequence of $\{\sigma_{k}\}$ with $\sigma_{k}\uparrow+\infty$, and set $\tau_{\min}\geq 4$ to ensure that $\sqrt{\tau_{\min}}-2\sqrt{\rho}>0$ for any $\rho\in[0,1)$. In contrast, as discussed in [58, Remark 2.1], ciPALM (which is also of relative-type but uses the error criterion (2.8)) should require $\rho<\frac{1}{3}$ to guarantee an asymptotic (super)linear convergence rate under the same error bound condition in Assumption A. This highlights another advantage of our ripALM, as it offers greater flexibility in choosing the tolerance parameter $\rho$.
Before closing this section, we summarize in Table 1 a brief comparison among the related inexact (p)ALMs, based on the discussions in Section 2 and the convergence results established in this section. From the comparison, we see that ripALM enjoys stronger theoretical guarantees than ES-iALM and, unlike ciPALM, does not require correction steps.
Table 1: A brief comparison among ES-iALM, Snipal, ciPALM, and ripALM. The convergence results for each algorithm are presented in the context of solving problem (1.2), which is consistent with our setting. In the table, “Seq. Conv.”: sequential convergence; “Abs.”: absolute-type; “Rel.”: relative-type; ✔E: be convergent under error bound condition.
Methods Type
Vanilla
(p)ALM
Proximal
Term
Boundedness Seq. Conv.
Asymptotic
(super)linear Rate
(under Error Bound)
$\{\bm{y}^{k}\}$ $\{\bm{x}^{k}\}$ $\{\bm{y}^{k}\}$ $\{\bm{x}^{k}\}$
ES-iALM [1, 14, 61] Rel. $\{\bm{x}^{k}\}$
Snipal [24] Abs. $\{(\bm{y}^{k},\bm{x}^{k})\}$
ciPALM [58] Rel.
$\{(\bm{y}^{k},\bm{x}^{k})\}$
ripALM [this work] Rel. E $\{(\bm{y}^{k},\bm{x}^{k})\}$
4 Applications of ripALM
In this section, we present the application of the ripALM in Algorithm 1 to two important problems: the quadratically regularized optimal transport problem and the basis pursuit denoising problem.
4.1 Quadratically regularized optimal transport
As an important variant of the classical optimal transport problem, the quadratically regularized optimal transport (QROT) problem introduces a quadratic regularization (i.e., ridge regularization) term into the objective function. This renders the problem strongly convex while preserving the sparsity of the optimal transport plan. These favorable properties make QROT a compelling alternative to the popular entropically regularized optimal transport problem, especially in applications where the sparse structure is desirable; see, e.g., [5, 16, 31] for more details. Mathematically, the QROT problem is given as follows:
$\min\limits_{X}~~\frac{\lambda}{2}\|X\|_{F}^{2}+\langle C,\,X\rangle\quad\mathrm{s.t.}\quad X\bm{1}_{n}=\bm{\alpha},~~X^{\top}\bm{1}_{m}=\bm{\beta},~~X\geq 0,$ (4.1)